# Is JPEG AI going to change image forensics?

### v1

@@ Abstract
In this paper, we investigate the counter-forensic effects of the forthcoming JPEG AI standard based on neural image compression, focusing on two critical areas: deepfake image detection and image splicing localization. Neural image compression leverages advanced neural network algorithms to achieve higher compression rates while maintaining image quality. However, it introduces artifacts that closely resemble those generated by image synthesis techniques and image splicing pipelines, complicating the work of researchers when discriminating pristine from manipulated content. We comprehensively analyze JPEG AI’s counter-forensic effects through extensive experiments on several state-of-the-art detectors and datasets. Our results demonstrate that an increase in false alarms impairs the performance of leading forensic detectors when analyzing genuine content processed through JPEG AI. By exposing the vulnerabilities of the available forensic tools we aim to raise the urgent need for multimedia forensics researchers to include JPEG AI images in their experimental setups and develop robust forensic techniques to distinguish between neural compression artifacts and actual manipulations.
@@ 1 Introduction
JPEG AI is a new standard in image compression developed to harness Neural Networks (NNs)’ power [3, 16]. Set to be published as an International Standard in early 2025 [46, 4], JPEG AI aims not only to achieve superior rate-distortion performance for enhanced visual quality but also to enable faster coding speeds and provide the flexibility to optimize human viewing and machine analysis [16, 47].
However, the introduction of JPEG AI poses new challenges for image forensics [40]. JPEG AI, as a novel standard, introduces unique forensic footprints that require dedicated study to understand and exploit [19, 79]. Additionally, existing literature indicates that neural image compression schemes leave distinctive upsampling traces similar to those found in images generated by NNs, such as deepfakes [20]. Deepfake image detection is particularly active, with numerous contributions highlighting that NN-based image generation techniques leave artifacts due to upsampling operators [81, 35, 37, 27, 59]. Furthermore, it is well-documented that locally manipulated (i.e., spliced) images may exhibit artifacts related to upsampling [75], too.
@@
This paper investigates whether state-of-the-art detectors designed for deepfake image detection and image splicing localization can mistakenly identify pristine JPEG AI-compressed images as malicious content. To achieve our goal, we consider several state-of-the-art detectors for both deepfake image detection and image splicing localization. We first collect their test datasets and compress them using the official JPEG AI Reference Software [10]. Then, we evaluate if the detectors present a pronounced tendency to misclassify JPEG AI compressed samples as malicious, varying the compression ratio. We also examine the detector’s performance on images encoded with traditional JPEG compression, confirming that artifacts introduced by JPEG AI are notably more misleading than those from conventional image compression methods. Finally, we preliminary show that incorporating JPEG AI images into the training pipeline of detectors helps mitigate this phenomenon.
The main findings of our experimental campaign are reported in Fig. 1. Our results prove that deepfake image detectors tend to classify pristine images compressed with JPEG AI as synthetic samples (top row), while image splicing localization techniques risk being fooled by JPEG AI, as many false alarms arise in the estimated tampering masks (bottom row). All these effects prove to be stronger as the JPEG AI compression ratio increases.
In summary, the main contributions of our paper are: i) we are the first to investigate deepfake image detection and image splicing localization tasks focusing on the official implementation of JPEG AI; ii) despite the exceptional visual quality of JPEG AI images, we demonstrate that state-of-the-art detectors are misled by interpolation artifacts, leading to false alarms when analyzing pristine content compressed with JPEG AI; iii) we reveal that this effect is incomparable with standard compression formats like JPEG; iv) we introduce JPEG AI samples into training to reduce the effects of this phenomenon.
@@ 2 JPEG AI background
General framework. In 2019, the JPEG Standardization Committee launched the JPEG AI initiative to create the first image coding specifications based on end-to-end learning [44]. At its core, JPEG AI is designed to transform an image into a latent tensor, which is then compressed and transmitted to the receiver, allowing the decoder to either reconstruct the image or carry out additional operations directly in the compressed domain using the latent tensor [16].
The general architecture of JPEG AI is depicted in Fig. 2 [45]. The encoder consists of a NN-based analysis transform that computes the latent information of the input image . Next, is processed by a latent domain prediction module, generating a latent residual , and by a hyperprior encoder, producing a hyperprior , as proposed in the foundational work by Ballé et al. [17]. For entropy decoding, a latent prediction network retrieves , which is finally input into the synthesis transform at the decoder for reconstructing the image . The entire system is trained end-to-end by minimizing a rate-distortion loss function. By adjusting the weight of the rate term, different models can be trained to achieve variable-rate compression, corresponding to various Bits Per Pixels (BPPs).
As a standardized format, JPEG AI enforces strict compliance requirements on manufacturers, limiting the customization options within the compression pipeline. In particular, the blue modules in Fig. 2 are formally normative. Although certain architectural modifications are permitted (indicated by the red modules), such as adjustments to the analysis transform, to the hyperprior encoder network, or retraining for specific content types to improve performance, these changes are restricted. The modified latent space must still function within the official JPEG AI framework to ensure compatibility [45].
This adherence to a fixed architecture means that images decoded with JPEG AI, even when produced by different manufacturers, are likely to display similar characteristics and similar forensic artifacts. As a consequence, studying the impact of JPEG AI on image forensics becomes a crucial area of investigation that warrants thorough exploration.
JPEG AI compression artifacts. Recently, it has been shown that neural compression codecs can leave frequency artifacts similar to those of synthetic generation techniques [19, 20]. These artifacts likely stem from the use of upsampling operators in the decoding process, and can be misleading for forensic detectors that rely on similar forensic footprints to accomplish their task. In this vein, we perform a similar analysis for the JPEG AI compression, following a well-known forensic procedure: we compute the average Fourier spectrum of noise residuals extracted from images under analysis [28, 59].
@@ 3 Case studies
In this paper, we aim to verify if forensic detectors fail in correctly working when dealing with JPEG AI-compressed images. We focus mainly on two forensics tasks: i) deepfake image detection and ii) image splicing localization. In the following, we illustrate each task and motivate our proposed experiments.
@@
Our study investigates whether state-of-the-art detectors produce high scores when processing genuine images compressed with JPEG AI. We aim to check if the artifacts introduced by JPEG AI (like the ones previously shown in Sec. 2) are misinterpreted by these detectors, leading to misclassifications of pristine samples. To do so, we compare the distributions of the scores obtained from pristine images and their JPEG AI-compressed versions.
Image splicing localization. With image splicing localization, we refer to the pixel-level identification of local manipulations in an image. For instance, this general manipulation manifests in the so-called “Photoshopped” or “cheapfake” [66] images, i.e., images where a malicious actor altered some pixels by copying content from another picture. To assess the integrity of the image under analysis, state-of-the-art detectors output an estimated real-valued tampering mask , whose values indicate the likelihood of manipulation for each pixel. Similarly to deepfake image detection, imposing a threshold on the likelihood values generates a binary tampering mask that classifies each pixel as manipulated or genuine.
@@ 4.1 Deepfake image detection
Forensic detectors. To have an exhaustive picture of the effects of JPEG AI on deepfake image detection, we select several state-of-the-art detectors ranging from Convolutional Neural Networks (CNNs) to transformers-based architectures, trained with different paradigms on diverse datasets. In particular, we choose the detector proposed in [81], available in two variants, i.e., [81]-A and [81]-B; the detectors proposed in [37, 28, 65, 59]; and, finally, the two approaches presented in [29], i.e., [29]-A and [29]-B. All these tools are publicly available, with both test code and models’ weights 111Complete experimental details in the Supplementary Material Sec. 7.. For each detector, we rely on the original authors’ implementation to execute our experiments.
Test set. To avoid any possible bias due to data distribution mismatching, we test each detector on the same test set used by the original authors. Our experiments’ final set of images comprehends pristine samples from CelebA [57], COCO [56], FFHQ [50], Imagenet [32], LAION [73], LSUN [86], and RAISE [30] datasets. We process all these images through JPEG AI compression, following a specific paradigm that we detail in Sec. 4.3.
Evaluation metrics. To quantify the counter-forensic effects of JPEG AI, we numerically compare the scores’ distributions of all detectors for all datasets of the genuine samples against their JPEG AI-compressed versions. We recall that each detector outputs a logit score that is thresholded at to distinguish between real () and synthetic () images. Our goal is to quantify how much JPEG AI compression is shifting to the right (i.e., above the threshold) the distribution of pristine samples. For this reason, we compare the two distributions (i.e., before and after the compression), considering these metrics:
Earth’s Mover Distance (EMD) [72]: a measure quantifying the “amount of work” to transform one distribution into another one by moving its mass. In our case, in the absence of JPEG AI counter-forensic effects, the EMD between pristine and JPEG AI samples should be ideally . Generally, the higher the EMD, the stronger the JPEG AI counter-forensic effect.
Area Under the Curve (AUC) and Balanced Accuracy (BA): two measures assessing the binary classification performance given two distributions. If the pristine and JPEG AI distributions are well-matched, these measures should be ideally ; otherwise, we expect higher values, approaching in case of a strong mismatch.
True Positive Rate (TPR): a measure quantifying how many JPEG AI samples could be potentially mistaken for deepfakes after thresholding their likelihood scores . Notice that such a metric requires ground-truth labels for each image. Ideally, both pristine and JPEG AI images should be labeled as “negative”. However, since we are interested in measuring the false alarms introduced by JPEG AI, we assign JPEG AI samples the “positive” ground truth. Thus, TPR represents the proportion of JPEG AI samples misclassified as synthetic, measuring how much JPEG AI confuses the classifier. Without counter-forensic effects, we expect the TPR to be close to ; conversely, strong counter-forensic effects can drive the TPR closer to , indicating high misclassification.
@@ 4.2 Image splicing localization
Forensic detectors. We focus on the latest and top-performing state-of-the-art techniques based on deep learning models. In particular, we choose two detectors: i) TruFor [38], a method based on a denoising CNN, and ii) MMFusion [78], a framework for the multi-modal fusion of different image splicing localization technique. Code and model weights are available online, and we rely on the original author’s implementation for our experiments.
Test set. We consider standard benchmarks in the field, i.e., CASIA [34], Columbia [42], Coverage [84], COCOGlide [38], and DSO1 [31, 25]. Both detectors were originally tested with excellent performances on these datasets, which comprise uncompressed samples (DSO1, Columbia, Coverage, COCOGlide) and JPEG images (CASIA). We compress all these images through JPEG AI by following the specifications provided in Sec. 4.3.
Evaluation metrics. We evaluate the performances of both detectors on the JPEG AI version of the datasets by comparing the ground truth mask against the estimated real-valued mask for each sample. Without counter-forensic effects, we expect a correct separation between pristine and tampered with pixels. On the contrary, in the case of the counter-forensic effect of JPEG AI, we expect a potential performance drop since JPEG AI compression might hinder the artifacts left by the local manipulation, and the presence of JPEG AI upsampling artifacts might cause false alarms even for pristine pixels. We consider the following metrics computed at a pixel level for each image:
and : we use the subscript S (i.e., splicing) to distinguish both metrics from the image-level metrics of the previous case study. Without counter-forensic effects, we expect and values around , i.e., a correct division between pristine and manipulated pixels. In the case of a strong counter-forensic effect of JPEG AI, the distributions of pristine and manipulated pixel values in risk being almost super-imposed to one another, achieving and values .
False Positive Rate (FPR): a measure indicating the probability of detecting a false alarm. In our case, it returns the probability that pristine pixels are mistakenly identified as manipulated after JPEG AI compression. To classify pristine versus tampered with pixels, the likelihood scores in are thresholded by the common threshold of . Without counter-forensic effects, we expect FPR to be close to ; contrarily, in the case of strong effects, FPR can approach , meaning heavy misclassification.
F1-score: the harmonic mean of precision and recall. In case of optimal localization performance, we expect F1-score to approach , decreasing its value as the counter-forensic effect increases until reaching its minimum in .
@@ 4.3 Dataset compression
We compressed the datasets illustrated above using the official JPEG AI Reference Software [10]. In particular, we used version 7.0 of the JPEG AI Verification Model [10] with the high operation point (HOP) configuration and enabled all tools to achieve the best coding performance [10]. We rely on six different BPP values, i.e., , and . We selected these values from the official target rates of the JPEG AI Call for Evidence Challenge [1]. With this approach, we have a believable overview of how the compression ratio affects the presence of artifacts and, in turn, the performance of forensic detectors.
@@ 5.1 Deepfake image detection
Preliminary analysis. Before analyzing the counter-forensic effect of JPEG AI, we first evaluate the various detectors on pristine and synthetic images not compressed with JPEG AI. This approach allows us to verify if detectors achieve acceptable performances for the standard real-vs-deepfake task. To do this, we include for every pristine test set its deepfake counterpart, selected by the original detector authors to depict similar content to the pristine one 1.
In this scenario, pristine samples represent the “negative” class, while deepfake samples are the “positive” class. It is worth noticing that the evaluation metrics presented in Sec. 4.1 for describing the similarity between pristine and JPEG AI samples are still valid for evaluating deepfake detection performances, although assuming an opposite role: a good deepfake detector should report high EMD values, with AUC and TPR both approaching the . On the contrary, for achieving good performances on JPEG AI images (i.e., resulting in a reduced JPEG AI counter-forensic effect), all these metrics should be as low as possible. For the sake of clarity, in Tab. 1, we visually summarize this double meaning associated with the employed metrics according to the specific scenario under investigation.
| EMD | AUC | BA | TPR | |
| Expected range | ||||
| Real-vs-synthetic | ||||
| Real-vs-JPEG AI |
@@
All detectors show satisfactory performances 222Complete results for all detectors and datasets available in Sec. 8 and Sec. 9 of the Supplementary Material., though the evaluation metrics are not uniform across all datasets. Given our goal of studying the effect of JPEG AI on misleading deepfake image detectors when processing genuine content, in the following, we discard results from datasets where detectors’ performances do not exceed an AUC of . Indeed, we want to analyze the counter-forensic effects of JPEG AI only when the initial dataset (before JPEG AI compression) is classified with acceptable performances by the detectors.
| Synthetic | BPP = 0.12 | BPP = 0.25 | BPP = 0.5 | BPP = 0.75 | BPP = 1.0 | |||||||||||||
| EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | |||||||||||||
| [81]-A | / | / | / | / | / | / | 9.73 / | 0.76 / | 0.21 | / | / | / | / | |||||
| [81]-B | / | / | / | / | / | / | 6.52 / | 0.68 / | 0.02 | / | / | / | / | |||||
| [37] | / | / | / | / | / | / | 11.31 / | 0.87 / | 0.29 | / | / | / | / | |||||
| [28] | / | / | / | / | / | / | 8.17 / | 0.79 / | 0.03 | / | / | / | / | |||||
| [65] | / | / | / | / | / | / | 2.33 / | 0.69 / | 0.09 | / | / | / | / | |||||
| [29]-A | / | / | / | / | / | / | 0.66 / | 0.72 / | 0.64 | / | / | / | / | |||||
| [29]-B | / | / | / | / | / | / | 0.77 / | 0.70 / | 0.54 | / | / | / | / | |||||
| [59] | / | / | / | / | / | / | 2.62 / | 0.83 / | 0.63 | / | / | / | / | |||||
Impact of compression ratio. We begin our analysis by illustrating the effect of JPEG AI on the score distributions obtained by the various detectors. We recall that, in this experiment, we test only pristine images and their JPEG AI versions. As previously conjectured, JPEG AI causes a right shift to all distributions, with the concrete risk of introducing false alarms when detecting deepfakes. As an example of this phenomenon, Figure 4 depicts the logit scores distribution of different detectors when processing original and JPEG AI samples from various datasets 2. Notice that, as the compression ratio increases (i.e., when we lower the BPPs), the distribution of the scores moves its center of gravity to the right, meaning that real samples can be potentially mistaken for deepfakes. In some cases, this effect is more progressive (see the first two plots), while in others, it is evident even at higher BPPs (last plot). Interestingly, a left shift of the distribution can also happen (second plot), though for very high BPPs only, where we observed that compression artifacts can be strongly reduced.
Tab. 2 reports the general results obtained by testing all datasets for each detector333We omit the BA results and BPP scenario for clarity. Complete results are again available in Sec. 8.. As a reference, we also report the detectors’ overall performance in the real-vs-synthetic scenario (first column). For extremely low BPP, all detection metrics resemble those achieved in the real-vs-synthetic setup. In the case of [37, 28, 59] detectors, JPEG AI compressed images sometimes carry even stronger “deepfake”-like artifacts than actual synthetic images. The high TPR, evaluated at threshold as it is commonly done in deepfake detection, also confirms that JPEG AI scores distribution moves above the threshold. All these elements tend to confirm that JPEG AI versions of pristine samples in low-rate settings can be mistaken for deepfake images. As expected, when increasing the BPP, JPEG AI images tend to align with the original ones. Nonetheless, the AUC is still around even when the JPEG AI compression ratio gets lower (see the case in which BPP is equal to ).
@@
JPEG AI image quality analysis. Given the performance drop of all detectors when tested against JPEG AI images, we now investigate if this loss might be due to the poor visual quality of JPEG AI compressed images with respect to the original ones. To do this, we compute the Multi-scale SSIM (MS-SSIM) and the Visual Information Fidelity (VIF) for all the JPEG AI images in the datasets when compared to their pristine versions. The JPEG committee proposes both metrics to evaluate the quality of JPEG AI implementations [1, 3] 444Additional results available in Sec. 8..
@@
| Synthetic | BPP = 0.12 | BPP = 0.25 | BPP = 0.50 | BPP = 0.75 | BPP = 1.0 | |
| EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | EMD / AUC / TPR | |
| Before | / / | / / | / / | / / | / / | / / |
| Retraining | / / | / / | / / | / / | / / | / / |
| Fine-tuning | / / | / / | / / | / / | / / | / / |
@@
To investigate this, we compare pristine images against their standard JPEG compressed versions using different Quality Factors (QFs). For example, Fig. 5 shows the distribution of scores for the detector presented in [65] over the Imagenet dataset. Notice that, differently from the JPEG AI case, the JPEG compressed images closely overlap with the original samples, meaning the absence of false alarms (scores’ distributions are all centered below the ) 2.
Finally, Fig. 6 depicts the average detection metrics of all detectors for all test pristine datasets in the two scenarios of pristine vs. JPEG AI (first column) and pristine vs. JPEG compression (second column). As we can quickly inspect, JPEG compression does not have the same effect as JPEG AI regarding “moving” the distribution of the scores above the threshold. In this case, pristine and JPEG compressed samples are both classified as real, i.e., their distributions almost overlap one with the other, regardless of the QF used.
@@
Introducing JPEG AI into training. Given that all the investigated detectors have not been trained with JPEG AI images, we conduct a last experiment by retraining the forensic detector recently proposed in [59] with JPEG AI samples. We consider the same training data used in the training phase of this detector, however we compress the of training and validation data with different BPPs (those considered in our previous analysis), equally distributed among the considered images. Notice that, in this case, we compress with JPEG AI both real and deepfake data. Then, we retrain this detector in two scenarios: (i) complete retraining over the same huge dataset used in [59], with of its samples compressed with JPEG AI; (ii) a fine-tuning of the already trained detector by exploiting only JPEG AI samples. In both cases, we follow the original training procedure [59].

### v2

@@ Abstract
In this paper, we investigate the counter-forensic effects of the new JPEG AI standard based on neural image compression, focusing on two critical areas: deepfake image detection and image splicing localization. Neural image compression leverages advanced neural network algorithms to achieve higher compression rates while maintaining image quality. However, it introduces artifacts that closely resemble those generated by image synthesis techniques and image splicing pipelines, complicating the work of researchers when discriminating pristine from manipulated content. We comprehensively analyze JPEG AI’s counter-forensic effects through extensive experiments on several state-of-the-art detectors and datasets. Our results demonstrate a reduction in the performance of leading forensic detectors when analyzing content processed through JPEG AI. By exposing the vulnerabilities of the available forensic tools we aim to raise the urgent need for multimedia forensics researchers to include JPEG AI images in their experimental setups and develop robust forensic techniques to distinguish between neural compression artifacts and actual manipulations.
@@ 1 Introduction
JPEG AI is a new standard in image compression developed to harness Neural Networks (NNs)’ power [3, 17]. Published as an International Standard on February 2025 [16], JPEG AI aims not only to achieve superior rate-distortion performance for enhanced visual quality but also to enable faster coding speeds and provide the flexibility to optimize human viewing and machine analysis [17, 51].
However, the introduction of JPEG AI poses new challenges for image forensics [45]. JPEG AI, as a novel standard, introduces unique forensic footprints that require dedicated study to understand and exploit [22, 86]. Additionally, existing literature indicates that neural image compression schemes leave distinctive upsampling traces similar to those found in images generated by NNs, such as deepfakes [23]. Deepfake image detection is particularly active, with numerous contributions highlighting that NN-based image generation techniques leave artifacts due to upsampling operators [88, 40, 42, 31, 63, 81]. Furthermore, it is well-documented that locally manipulated (i.e., spliced) images may exhibit artifacts related to upsampling [80], too.
@@
This paper investigates whether state-of-the-art algorithms designed for deepfake image detection and image splicing localization can mistakenly identify JPEG AI-compressed images as malicious content. To achieve our goal, we consider several state-of-the-art detectors from both tasks. We first collect their test datasets and compress them using the official JPEG AI Reference Software [10]. Then, we evaluate if the detectors present a pronounced tendency to misclassify JPEG AI compressed samples as malicious, varying the compression ratio. We also examine the detector’s performance on images encoded with traditional JPEG compression, confirming that artifacts introduced by JPEG AI make detectors behave in a sort of opposite way to conventional compression methods. Finally, we preliminary show that incorporating JPEG AI images into the training pipeline of detectors may help mitigate this phenomenon, and also experiment with double JPEG AI compression.
The main findings of our experimental campaign are reported in Fig. 1. We prove that deepfake image detectors tend to classify images compressed with JPEG AI as synthetic samples (top row), while image splicing localization techniques risk being fooled by JPEG AI, as many missed detections and false alarms may arise in the estimated tampering masks (bottom row). All these effects prove to be stronger as the JPEG AI compression ratio increases.
In summary, the main contributions of our paper are: i) we are the first to investigate deepfake image detection and image splicing localization tasks focusing on the official implementation of JPEG AI; ii) despite the exceptional visual quality of JPEG AI images, we demonstrate that state-of-the-art detectors are misled by their interpolation artifacts; iii) we reveal that this effect is incomparable with standard compression formats like JPEG; iv) we introduce JPEG AI samples into training to reduce the effects of this phenomenon; v) we investigate the counter-forensic effects of double JPEG AI compression.
@@ 2 JPEG AI background
General framework. In February 2025, the JPEG Standardization Committee published JPEG AI as the first international image coding standard based on end-to-end learning [16]. The general architecture of JPEG AI is depicted in Fig. 2 [50]. The basic functioning follows the same of many other neural image compression algorithms: it first transforms an image into a latent tensor, which is then compressed and transmitted to the receiver, allowing the decoder to either reconstruct the image or carry out additional operations directly in the compressed domain exploiting the latent tensor [17]. The entire JPEG AI codec is trained end-to-end by minimizing a rate-distortion loss function. By adjusting the weight of the rate term, different models can be trained to achieve variable-rate compression, corresponding to various Bits Per Pixel (BPP). We refer the reader to the supplemental material, Sec. 7, for more details on each architectural module.
While many neural image compression techniques exist [67] (please refer to the supplemental material for more discussion), in our analysis we specifically focus on JPEG AI given its forthcoming impact. Indeed, JPEG AI presents interesting characteristics that are likely to spread its adoption among users, including support for a wide array of devices with different characteristics [16] and the direct involvement of technology companies [9].
It is noteworthy to mention that JPEG AI, as a standardized format, enforces strict compliance requirements on manufacturers, limiting customization options within the compression pipeline. In particular, the blue modules in Fig. 2, among which we can notice the decoder, are formally normative. This adherence to a fixed architecture suggests that images decoded with JPEG AI, even when produced by different manufacturers, will likely display similar characteristics and forensic artifacts.
JPEG AI compression artifacts. Recently, it has been shown that neural compression codecs can leave frequency artifacts similar to those of synthetic generation techniques [22, 23]. These artifacts presumably stem from the use of upsampling operators in the decoding process, and can be misleading for forensic detectors that rely on similar forensic footprints to accomplish their task. In this vein, we perform a similar analysis for JPEG AI compression, following a well-known forensic procedure, i.e., we compute the average Fourier spectrum of noise residuals extracted from images under analysis [32, 63].
@@ 3 Case studies
In this paper, we aim to verify if forensic detectors fail to work correctly when dealing with JPEG AI-compressed images. We focus mainly on two forensics tasks: i) deepfake image detection and ii) image splicing localization. In the following, we illustrate each of them and motivate our proposed experiments.
@@
Our study investigates whether state-of-the-art detectors produce high scores when processing images compressed with JPEG AI. We aim to check if the artifacts introduced by JPEG AI (like the ones previously shown in Sec. 2) are misinterpreted by these detectors leading to the misclassification of pristine samples as deepfakes or to alterations of the scores of deepfake images. To do so, we compare the distributions of the scores obtained from JPEG AI-uncompressed images and their JPEG AI-compressed versions.
Image splicing localization. With image splicing localization, we refer to the pixel-level identification of local manipulations in an image. For instance, this general manipulation manifests in the so-called “Photoshopped” or “cheapfake” [71] pictures, i.e., images where a malicious actor altered some pixels by copying content from another picture. To assess the integrity of the image under analysis, state-of-the-art detectors output an estimated real-valued tampering mask , whose values indicate the likelihood of manipulation for each pixel. Similarly to deepfake image detection, imposing a threshold on the likelihood values generates a binary tampering mask that classifies each pixel as manipulated or genuine.
@@ 4.1 Deepfake image detection
Forensic detectors. To have an exhaustive picture of the effects of JPEG AI on deepfake image detection, we select several state-of-the-art detectors ranging from Convolutional Neural Networks (CNNs) to transformers-based architectures, trained with different paradigms on diverse datasets. In particular, we choose the detector proposed in [88], available in two variants, i.e., [88]-A and [88]-B, the detectors proposed in [42, 32, 70, 63, 81] and, finally, the two approaches presented in [34], i.e., [34]-A and [34]-B. All these tools are publicly available, with both test code and model weights. We rely on the original authors’ implementation for each detector to execute our experiments 111Complete experimental details in the Supplementary Material Sec. 8. All code for replicating our experiments is available here. .
Test set. To avoid any possible bias due to data distribution mismatching, we test each detector on the same test set used by the original authors. Our experiments’ final set of images comprehends pristine samples from CelebA [61], COCO [60], FFHQ [54], Imagenet [37], LAION [78], LSUN [95], and RAISE [35] datasets. For every pristine test set, we include a deepfake counterpart, selected by the original authors to depict similar content to the pristine one1. We process all these images through JPEG AI compression, following a specific paradigm we detail in Sec. 4.3.
Evaluation metrics. To quantify the counter-forensic effects of JPEG AI, we compare pristine and deepfake scores’ distributions considering two scenarios, i.e., before and after JPEG AI compression is applied to our test images. Our goal is to quantify how much JPEG AI alters these scores’ distribution.
As previously conjectured, we expect JPEG AI to move the score distribution of pristine samples to the “deepfake” area, i.e., above the threshold, potentially producing more false alarms. Moreover, we later show that JPEG AI compression does not only modify pristine samples’ distributions, but it can also affect those of synthetic images. To quantify these modifications, we use standard binary classification metrics, namely the Area Under the Curve (AUC), the Balanced Accuracy (BA), the False Positive Rate (FPR), and the False Negative Rate (FNR) evaluated at threshold. The FPR quantifies the ratio of pristine samples misclassified as synthetic (i.e., the false alarms), while the FNR represents the ratio of synthetic samples misclassified as pristine. If the JPEG AI has no effect, these metrics should remain the same before and after the compression.
@@ 4.2 Image splicing localization
Forensic detectors. We focus on the latest and top-performing state-of-the-art techniques based on deep learning models. In particular, we choose three detectors: i) TruFor [43], a method based on a denoising CNN; ii) MMFusion [85], a framework for the multi-modal fusion of different image splicing localization techniques; iii) ImageForensicsOSN (IFOSN) [92], a technique robust to social network compression schemes. Code and model weights are available online, and we rely on the original author’s implementation for our experiments.
Test set. We consider standard benchmarks in the field, i.e., CASIA [39], Columbia [47], Coverage [91], COCOGlide [43], and DSO-1 [36, 28]. All detectors were originally tested with excellent performances on these datasets, which comprise uncompressed samples (DSO-1, Columbia, Coverage, COCOGlide) and JPEG images (CASIA). We compress all these images through JPEG AI by following the specifications provided in Sec. 4.3.
Evaluation metrics. In this case as well, we evaluate the performances of all detectors on the JPEG AI-uncompressed version of the datasets and their JPEG AI-compressed version. In specific, we compare the ground truth mask against the estimated real-valued mask for each sample before and after the compression.
Without counter-forensic effects, we expect a correct separation between pristine and tampered with pixels. On the contrary, in the case of the counter-forensic effect of JPEG AI, we expect a potential performance drop since JPEG AI compression might hinder or introduce new artifacts. To measure this behavior, we consider standard metrics for the image splicing localization task, namely the pixel level AUC, the BA and the F1-score. The BA and the F1-score are evaluated at the threshold.
@@ 4.3 Dataset compression
We compress the datasets illustrated above using the official JPEG AI Reference Software [10]. In particular, we use version 7.0 of the JPEG AI Verification Model [10] with the High Operation Point (HOP) configuration and enable all tools to achieve the best coding performance [10]. We rely on six different BPP values, i.e., , and . We select these values from the official target rates of the JPEG AI Call for Evidence Challenge [1]. With this approach, we have a believable overview of how the compression ratio affects the presence of artifacts and, in turn, the performance of forensic detectors.
@@ 5.1 Deepfake image detection
Preliminary analysis. Before analyzing the counter-forensic effect of JPEG AI, we first evaluate the various detectors on pristine and synthetic images not compressed with JPEG AI. This approach allows us to verify if detectors achieve acceptable performances for the standard real-vs-deepfake task. The first column of Tab. 1 reports the results averaged across all datasets. On average, all detectors show satisfactory performances222Complete results for all detectors and all datasets are available in Sec. 9 and Sec. 10 of the Supplementary Material..
@@
Given our goal of studying the effect of JPEG AI on misleading deepfake image detectors, in the following, we discard results from datasets where detectors’ performances do not exceed an AUC of . Indeed, we want to analyze the counter-forensic effects of JPEG AI only when the initial dataset (before JPEG AI compression) is classified with acceptable performances by the detectors.
Impact of compression ratio. We begin our analysis by illustrating the effect of JPEG AI on the score distributions. Figure 4 depicts the logit scores distribution of detector [42] when processing original format and JPEG AI compressed samples from the COCO dataset2.
Interestingly, as the compression ratio increases (i.e., when we lower the BPPs), the distribution of the scores moves its center of gravity to the right (i.e., above the threshold), almost independently from the class of the samples. This behavior has different implications. For pristine samples, we expect an increase in false positives. This right shift may cause fewer false negatives for deepfake samples instead.
To confirm this hypothesis quantitatively, Tab. 1 (starting from the second left column) reports the average results obtained by testing each detector on all its datasets333We omit the AUC results and BPP scenario for clarity. Complete results are available in the Supplementary Material Sec. 9.. For extremely low BPPs, all detectors exhibit a drop in the BA, especially for [81, 32, 42, 63], and they all experience an increase in the FPR. All these elements tend to confirm that, in low-rate settings, the score distributions of JPEG AI pristine samples move above the threshold. Hence, these images can be mistaken for deepfakes. On the contrary, for all detectors except [88]-B and [70], the FNR decreases for aggressive compression settings. This phenomenon seems to corroborate our intuition: JPEG AI artifacts might “overwrite” synthetic generation traces at very low BPP values. However, most detectors still mistake these artifacts as deepfake generation clues, thus reducing the FNR and ending up improving the classification performances of synthetic images.
As might be expected, when increasing the BPP, JPEG AI score distributions tend to align more closely with the original ones. Nonetheless, the BA can still be less than , and the FPR risks to be high even when the JPEG AI compression ratio gets lower (see the case in which BPP is equal to for detectors [34]-A, [34]-B and [81]). Moreover, when the BPP increases, also a left-shift of the scores of synthetic samples can occur, thus resulting in a slight enhancement of FNR and reducing the detection performances (see, for instance, detectors [88]-A, [42], [32], [70] and [63]).
@@
JPEG AI image quality analysis. Given the performance drop of all detectors when tested against JPEG AI images, we now investigate if this loss might be due to the poor visual quality of JPEG AI compressed images with respect to the original ones. To do this, we compute the Multi-scale SSIM (MS-SSIM) and the Visual Information Fidelity (VIF) for all the JPEG AI images in the datasets, comparing them to their JPEG AI-uncompressed versions. The JPEG AI committee proposes both metrics to evaluate the quality of JPEG AI implementations [1, 3] 444Additional results available in Sec. 9..
@@
| Original format | BPP = | BPP = | BPP = | BPP = | BPP = | |||||||||||||
| Before | ||||||||||||||||||
| Retraining | ||||||||||||||||||
| Fine-tuning | ||||||||||||||||||
@@
To investigate this, we compress our datasets using different Quality Factors (QFs), then compare the distribution of the scores before and after JPEG compression. Fig. 5 shows an example, i.e., the score distribution for the detector [34]-A from the LSUN dataset2. Notice that, differently from the JPEG AI case, the scores distribution of pristine JPEG-compressed images closely overlaps with those obtained from JPEG-uncompressed samples, meaning the absence of false alarms (both distributions are all centered below the ). On the other hand, when compressing deepfake images, the scores distribution moves below the as we increase the QF, meaning a significant increase in the number of false negatives.
To confirm this observation, Tab. 3 reports the average classification metrics for all datasets and detectors before (first left column) and after (starting from the second left column) JPEG varying the QF used in the compression. In this case, the BA has a general drop, which is more evident in the case of lower QFs. However, unlike the JPEG AI scenario, the FPR stays constant or even lowers, indicating that the scores distribution of pristine samples remains or moves even below the threshold. On the other hand, JPEG compression, apart from detector [32], consistently affects the FNR, which is generally higher, especially for [88]-A, [42], [34]-A, [81] and [63]. This suggests that JPEG compression shifts the scores distribution of deepfake images below the , effectively “canceling” out forensic artifacts that are crucial for deepfake detection. This behavior is diametrically opposed to that of JPEG AI, which essentially affects the detection of pristine content, increasing the number of false positives.
@@
Introducing JPEG AI into training. Given that all the investigated detectors have not been trained with JPEG AI images, we conduct a last experiment by retraining the forensic detector recently proposed in [63] with JPEG AI samples. We consider the same data used in the training phase of this detector. However, we compress the of training and validation samples with the BPP values considered in our previous analysis, equally distributed among the considered images. Then, we retrain this detector in two scenarios: (i) complete retraining over the same dataset used in [63], with of its samples compressed with JPEG AI; (ii) a fine-tuning of the already trained detector by exploiting only JPEG AI samples. In both cases, we follow the original training procedure [63].